What is the probability that in two rolls of a single die, a man will roll a 2 on the first roll and a number greater than five on the second ?
Answer by ewatrrr(24785)


Hi
two rolls of a single die (6 faces)
2 on the first roll AND a number greater than five on the second (6 the only choice)
P = (1/6)(1/6) = 1/36
Wish You the Best in your Studies.